A plane flying with the wind flew 225 miles in 2 hours. It returned to its origin flying against the wind and took 3 hours 45 minutes to cover the same distance. How fast was the wind blowing? I came up with the following, did I do this correctly? Thanks for your help
V= Air Speed
W= Wind Speed
2(V + W)=225
2V + 2W =225
2V = 225
112.5 is air speed
3.75(112.5 - W)=225
421.88 - 3.75W = 225
-3.75W = -196.88
W = 52.50 mph

Answer by Paul(988)

Let the speedd of the plane be x
let the speed of the wind be y
2(x+y)=225 --> solve for y
2x+2y-225
2y=225-2x
y=112.5-x (subsitution)

3h and 45 minutes = 3.75hours.

3.75x-3.75y=225
3.75-3.75(112.5-x)=225
3.75-421.875+3.75=225
7.5x=646.875
x=86.25
y=112.5-86.25
y=26.25

Hence, the speed of the plane is 86/25mph, and the speed of the wind is 26.25mph.
Paul.